# Empowering LLMs to Understand and Generate Complex Vector Graphics

### v1

@@ 3 Vector Instruction Following Data
[…]
@@ 4.1 SVG Semantic Tokens
For an input SVG $\mathbf{X}_{v}$, we convert it from raw code into a structured representation. To accomplish this, we defined 55 SVG semantic tokens $s_{i}\in\mathbb{R}^{55}$ (including 15 tag tokens, 30 attribute tokens, 10 path command tokens). These SVG tokens are used to replace all tags and attributes in the SVG source code, thus preventing the textual encoding of SVG tags and attributes as regular text. For example, the tag <path> will be tokenized as an SVG semantic token, rather than the literal “path” by the tokenizer. This ensures the uniqueness of SVG tags and attributes, and allows for their efficient integration into LLMs in a manner that is consistent with SVG definitions and optimizes token initialization. We adapt the embedding layer $\mathbf{W}{\mathrm{emb}}\in\mathbb{R}^{|\mathcal{V}|^{\prime}}$ to learn the embeddings of these new tokens, where $|\mathcal{V}|^{\prime}:=|\mathcal{V}|+55$ represents the sum of the size of the original vocabulary and the additional SVG tokens. Each new token is initialized based on the semantic average of its descriptive text, as defined by the equation:
@@ 4.3 Training
where $\theta$ represents the trainable parameters of the model, $\mathbf{X}_{\mathrm{inst}}$ and $\mathbf{X}_{\mathrm{a}}$ are the tokens corresponding to the instructions and answers for all preceding turns before the current prediction token $\bm{x}_{i}$, respectively. The term $\hat{\bm{x}}_{i-1}=\bm{x}_{i-1},\cdots,\bm{x}_{1}$ denotes the sequence of tokens that have been predicted in previous steps.
@@ 5.3 Ablation Study
Analysis of the LLM4SVG Architecture. The performance of LLM4SVG is significantly influenced by the underlying base model, especially concerning how the LLM tokenizer manages continuous numeric tokens. For example, in the Qwen models [59, 52], during the tokenization process, all numbers and decimal points are converted into a single token. As a result, Qwen is unable to properly handle continuous numerical data, leading to poor performance in tasks such as coordinate and hexadecimal color prediction, making it difficult to generate complete and coherent SVGs. LLMs like GPT-2 [39], which enumerate all numbers within 1000 as individual tokens in the tokenizer, exhibit a stronger understanding of numerical data. This capability allows such models to generate more complete and coherent SVGs, leading to a more visually appealing and harmonious output.
Analysis of SVG Semantic Tokens. During the model training process, we expanded the tokenizer by incorporating 55 SVG-specific semantic tokens. This modification enables SVG tags and attributes to be identified with distinct tokens, thereby differentiating them from general text. For instance, prior to this addition, the word path could ambiguously refer to an SVG tag or merely a physical path. However, with the introduction of a dedicated path token, the model is now equipped to clearly differentiate between the SVG path tag and a literal path. This distinction aids the model in more effectively and accurately learning the SVG generation template, enhancing its ability to produce precise and contextually appropriate SVGs.
@@ 6 Conclusion & Discussion
  • In Section D, we present additional qualitative results of our LLM4SVG, showcasing its ability to generate SVGs with high editability and superior visual quality.
@@ Appendix A Comparison with LLM-based Methods
As illustrated in Fig.S3, it is evident that our proposed LLM4SVG outperforms other methods in terms of overall visual effect and detail expression. Specifically, most LLMs struggle to generate complete SVG images, for example, butterfly, or produce outputs that accurately align with the provided textual descriptions, such as two-hump camel illustrated in the last third example. Some recent LLMs perform relatively better, such as GPT-4o[2], GPT-o1-preview [2], and Claude 3.5-sonnet [35]. However, their results are still less satisfactory as the shapes are overly simplistic (e.g., flute) and the colors lack harmony (e.g., a bed). In contrast, the results of our LLM4SVG are more complete, with shapes that are more diverse, colors that are more harmonious, and semantics that are more closely aligned with the prompts.
@@
Figure S4: Example SVG Code Generated by GPT-o1-preview [2], Claude 3.5-sonnet [35], and Our LLM4SVG.
@@
Appendix B Dataset and Proprecessing Pipeline
@@ Appendix C More Details about User Study
Figure S6: Screenshot of Our Questionnaires.
Figure S7: Additional Results Generated by Our LLM4SVG.
We conducted a user survey to evaluate the effectiveness and practicality of the SVGs generated by our method LLM4SVG and two other popular LLMs, GPT4-o and Claude-3.5. Specifically, the user study was structured as follows:
@@
    Result Calculation: This user study involves 37 volunteers from backgrounds in computer science and the arts. Each volunteer was required to complete between 1 and 3 questionnaires. Scores presented in Table 3 of our manuscript were calculated by averaging the ratings for SVGs within the same category. For “Prompt Alignment” and “Visual Quality”, the ratings were adjusted by a coefficient $\alpha=0.2$, such that a score of 1 translates 0.2, and a score of 5 counts to 1. For “Pick Score”, a “Yes” was scored as 1, while a “No” was scored as 0.
Appendix D Additional Qualitative Results

### v2

@@ 3 Vector Instruction Following Data
Figure 2: An Overview of LLM4SVG. Our LLM4SVG is capable of understanding and generating SVGs effectively. (1) During the training phase, we provide both the original SVG code $\mathbf{X}_{v}$ and the corresponding instruction data $\mathbf{X}_{\mathrm{inst}}$ as input. For the understanding task, we use detailed descriptions $\mathbf{X}_{a}$ generated by GPT-4 [2] as the training labels. For the generation task, the SVG code portion is masked and serves as the target that the model needs to predict. (2) During the inference phase, for the understanding task, given an SVG source code, the model generates a description that aligns with the semantics expressed by the SVG. For the generation task, the model generates an SVG based on the input text prompt. During both training and inference phases, the rendered image $\mathbf{X}_{\mathrm{img}}$ of the SVG can be used as conditional input to the model, guiding the content that the model understands or generates.
@@ 4.1 SVG Semantic Tokens
For an input SVG $\mathbf{X}_{v}$, we convert it from raw code into a structured representation. To accomplish this, we defined 55 SVG semantic tokens $s_{i}\in\mathbb{R}^{55}$ (including 15 tag tokens, 30 attribute tokens, 10 path command tokens, details are provided in Sec. D and Fig.  S1 of Supplementary). These SVG tokens are used to replace all tags and attributes in the SVG source code, thus preventing the textual encoding of SVG tags and attributes as regular text. For example, the tag <path> will be tokenized as an SVG semantic token, rather than the literal “path” by the tokenizer. This ensures the uniqueness of SVG tags and attributes, and allows for their efficient integration into LLMs in a manner that is consistent with SVG definitions and optimizes token initialization. We adapt the embedding layer $\mathbf{W}_{\mathrm{emb}}\in\mathbb{R}^{|\mathcal{V}|^{\prime}}$ to learn the embeddings of these new tokens, where $|\mathcal{V}|^{\prime}:=|\mathcal{V}|+55$ represents the sum of the size of the original vocabulary and the additional SVG tokens. Each new token is initialized based on the semantic average of its descriptive text, as defined by the equation:
@@ 4.3 Training
where $\theta$ represents the trainable parameters of the model, $\mathbf{X}_{\mathrm{inst}}$ and $\mathbf{X}_{\mathrm{a}}$ are the tokens corresponding to the instructions and answers for all preceding turns before the current prediction token $\bm{x}_{i}$, respectively. The term $\hat{\bm{x}}_{i-1}\coloneqq(\bm{x}_{i-1},\cdots,\bm{x}_{1})$ denotes the sequence of tokens that have been predicted in previous steps.
@@ 5.3 Ablation Study
Analysis of the LLM4SVG Architecture. The performance of LLM4SVG is significantly influenced by the underlying base model, especially concerning how the LLM tokenizer manages continuous numeric tokens. For example, in the Qwen models [59, 52], during the tokenization process, all numbers and decimal points are converted into a single token. As a result, Qwen is unable to properly handle continuous numerical data, leading to poor performance in tasks such as coordinate and hexadecimal color prediction, making it difficult to generate complete and coherent SVGs. LLMs like GPT-2 [39], which enumerate all numbers within 1,000 as individual tokens in the tokenizer, exhibit a stronger understanding of numerical data. This capability allows such models to generate more complete and coherent SVGs, leading to a more visually appealing and harmonious output.
Analysis of SVG Semantic Tokens. During the model training process, we expanded the tokenizer by incorporating 55 SVG-specific semantic tokens. This modification enables SVG tags and attributes to be identified with distinct tokens, thereby differentiating them from general text. For instance, prior to this addition, the word “path” could ambiguously refer to an SVG tag or merely a physical path. However, with the introduction of a dedicated path token, the model is now equipped to clearly differentiate between the SVG path tag and a literal path. This distinction aids the model in more effectively and accurately learning the SVG generation template, enhancing its ability to produce precise and contextually appropriate SVGs.
@@ 6 Conclusion & Discussion
  • In Section D, we provide the details of the SVG semantic tokens.
  • In Section E, we present additional qualitative results of our LLM4SVG, showcasing its ability to generate SVGs with high editability and superior visual quality.
@@ Appendix A Comparison with LLM-based Methods
As illustrated in Fig.S3, it is evident that our proposed LLM4SVG outperforms other methods in terms of overall visual effect and detail expression. Specifically, most LLMs struggle to generate complete SVG images, for example, butterfly, or produce outputs that accurately align with the provided textual descriptions, such as two-hump camel illustrated in the last third example. Some recent LLMs perform relatively better, such as GPT-4o, GPT-o1-preview [2], and Claude 3.5-sonnet [35]. However, their results are still less satisfactory as the shapes are overly simplistic (e.g., flute) and the colors lack harmony (e.g., a bed). In contrast, the results of our LLM4SVG are more complete, with shapes that are more diverse, colors that are more harmonious, and semantics that are more closely aligned with the prompts.
@@
Figure S4: Example SVG Code Comparison. SVG Code Generated by GPT-o1-preview [2], Claude 3.5-sonnet [35], and Our LLM4SVG.
@@
Appendix B Dataset and Preprocessing Pipeline
@@
Figure S6: Screenshot of Our Questionnaires.
Category Token Description
SVG Container Tags [<|START_OF_SVG|>] start of svg
[<|END_OF_SVG|>] end of svg
[<|start_of_g|>] start of svg group
[<|end_of_g|>] end of svg group
SVG Shape Tags [<|svg_path|>] svg path element
[<|svg_circle|>] svg circle element
[<|svg_rect|>] svg rectangle element
[<|svg_ellipse|>] svg ellipse element
[<|svg_polygon|>] svg polygon element
[<|svg_line|>] svg line element
[<|svg_polyline|>] svg polyline element
[<|svg_text|>] svg text element
SVG Gradient Tags [<|svg_linearGradient|>] svg linear gradient element
[<|svg_radialGradient|>] svg radial gradient element
[<|svg_stop|>] svg stop element
Path Commands [<|moveto|>] svg path command, move to
[<|lineto|>] svg path command, line to
[<|horizontal_lineto|>] svg path command, horizontal line to
[<|vertical_lineto|>] svg path command, vertical line to
[<|curveto|>] svg path command, curve to
[<|smooth_curveto|>] svg path command, smooth curve to
[<|quadratic_bezier_curve|>] svg path command, quadratic bezier curve
[<|smooth_quadratic_bezier_curveto|>] svg path command, smooth quadratic bezier curve
[<|elliptical_Arc|>] svg path command, elliptical arc
[<|close_the_path|>] svg path command, close the path, close-form
Attribute Tokens [<|id|>] svg element attribute id
[<|d|>] svg element attribute define the path
[<|fill|>] svg element attribute fill
[<|stroke-width|>] svg element attribute stroke-width
[<|stroke-linecap|>] svg element attribute stroke-linecap
[<|stroke|>] svg element attribute stroke
[<|opacity|>] svg element attribute opacity
[<|transform|>] svg element attribute transform
[<|gradientTransform|>] svg element attribute gradient transform
[<|offset|>] svg element attribute offset
[<|width|>] svg element attribute width
[<|height|>] svg element attribute height
[<|cx|>] svg element attribute x coordinate of circle center
[<|cy|>] svg element attribute y coordinate of circle center
[<|rx|>] svg element attribute x radius of ellipse
[<|ry|>] svg element attribute y radius of ellipse
[<|r|>] svg element attribute radius of circle
[<|points|>] svg element attribute points
[<|x1|>] svg element attribute x1 coordinate
[<|y1|>] svg element attribute y1 coordinate
[<|x2|>] svg element attribute x2 coordinate
[<|y2|>] svg element attribute y2 coordinate
[<|x|>] svg element attribute x coordinate
[<|y|>] svg element attribute y coordinate
[<|fr|>] svg element attribute fr
[<|fx|>] svg element attribute fx
[<|fy|>] svg element attribute fy
[<|href|>] svg element attribute href
[<|rotate|>] svg element attribute rotate
[<|font-size|>] svg element attribute font-size
Table S1: SVG Semantic Tokens Defined by Our LLM4SVG. We define 15 tag tokens (including 4 SVG container tags, 8 SVG geometry tags, and 3 SVG gradient tags), 30 attribute tokens, and 10 path command tokens in our LLM4SVG. The “Token” field corresponds to the Token defined in Table 1. The “Description” field is used to initialize the Token.
Figure S7: Additional Results Generated by Our LLM4SVG.
@@ Appendix C More Details about User Study
We conducted a user survey to evaluate the effectiveness and practicality of the SVGs generated by our method LLM4SVG and two other popular LLMs, GPT-4o [2] and Claude-3.5 [35]. Specifically, the user study was structured as follows:
@@
    Result Calculation: This user study involves 37 volunteers from backgrounds in computer science and the arts. Each volunteer was required to complete between 1 and 3 questionnaires. Scores presented in Table 3 of our manuscript were calculated by averaging the ratings for SVGs within the same category. For “Prompt Alignment” and “Visual Quality”, the ratings were adjusted by a coefficient $\alpha=0.2$, such that a score of 1 translates to 0.2, and a score of 5 counts to 1. For “Pick Score”, a “Yes” was scored as 1, while a “No” was scored as 0.
Appendix D SVG Semantic Tokens
For an input SVG $\mathbf{X}_{v}$, we convert it from raw code into a structured representation. As shown in Tab. S1, we present a detailed taxonomy of the SVG semantic tokens employed in LLM4SVG, including 15 tag tokens, 30 attribute tokens and 10 path command tokens. These SVG tokens are used to replace all tags and attributes in the SVG source code, thus preventing the textual encoding of SVG tags and attributes as regular text. This ensures the uniqueness of SVG tags and attributes, and allows for their efficient integration into LLMs in a manner that is consistent with SVG definitions. The “Description” field is utilized to initialize the SVG Tokens based on Equation 1.
Appendix E Additional Qualitative Results